# On a novel UCP result and its application to inverse conductive scattering

### v2

@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
$|\eta_{1}u_{2}^{+}(\mathbf{0})|=0.$
However, this contradicts to the admissible condition $u_{2}^{+}(\mathbf{0})\neq 0$ due to $\eta_{1}\neq 0$.
Case (b): $\partial_{1}u_{2}^{+}(\mathbf{0})\not=0$, $u_{2}^{+}(\mathbf{0})=\partial_{2}u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), we have the following integral equality:
$\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma=-\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma$
$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.18)
@@
$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})|}{\tau^{2}}\left|\frac{\cos\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\cos\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau}),$ (3.20)
@@
$|\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})|=0,$
which implies that $\partial_{1}u_{2}^{+}(\mathbf{0})=0$ by noting $\eta_{1}\not=0$. Hence, we obtain a contradiction.
Case (c): $\partial_{2}u_{2}^{+}(\mathbf{0})\not=0,$ $u_{2}^{+}(\mathbf{0})=\partial_{1}u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), it holds that
$\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma=-\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma$
$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.21)
@@
$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})|}{\tau^{2}}\left|\frac{\sin\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\sin\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau})$ (3.22)
@@
$|\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})|=0,$
Hence, it yields that $\partial_{2}u_{2}^{+}(\mathbf{0})=0$ since $\eta_{1}\neq 0$, which contradicts to $\partial_{2}u_{2}^{+}(\mathbf{0})\not=0$.
Case (d): $\partial_{1}u_{2}^{+}(\mathbf{0})\not=0$,$\partial_{2}u_{2}^{+}(\mathbf{0})\not=0$, $u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), we have
$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}$ $(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\sigma-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma$
$-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\psi(\mathbf{x})\mathrm{d}\sigma+I_{4}.$ (3.23)
@@
$\mathcal{N}=$ $(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{M})e^{2\mathrm{i}\vartheta_{m}}$
$\quad+(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{m})e^{2\mathrm{i}\vartheta_{M}},$ (3.24)
@@
$\mathcal{N}=$ $(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{M})e^{-2\mathrm{i}\vartheta_{m}}$
$\quad+(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{m})e^{-2\mathrm{i}\vartheta_{M}}.$ (3.25)
Let us consider the real and imaginary parts of $\partial_{1}u_{2}^{+}(\mathbf{0})$ and $\partial_{2}u_{2}^{+}(\mathbf{0})$, seperately. Denote $\partial_{1}u_{2}^{+}(\mathbf{0})=a_{1}+b_{1}\mathrm{i}$ and $\partial_{2}u_{2}^{+}(\mathbf{0})=a_{2}+b_{2}\mathrm{i}$, where $a_{i},b_{i}\in\mathbb{R}\ (i=1,2).$ For (3.24), it follows that
@@
Therefore, by (3.28), we have $a_{i}=b_{i}=0$ for $i=1,2$, which indicates that $\partial_{1}u_{2}^{+}(\mathbf{0})=\partial_{2}u_{2}^{+}(\mathbf{0})=0$. This leads to a contradiction.
@@
The coefficients of $q_{i,\ell}$ are real, where $q_{i,\ell}$ follows the form specified in (3.1), while $\tilde{\eta}_{\ell}$ is a real constant. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega;q_{\ell},\eta_{\ell})$. Assume that
@@
The coefficients of $q_{i,\ell}$ are real, where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ is a real constant. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega_{\ell};q_{\ell},\eta_{\ell})$. Assume that
@@
for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$. Then we have $N_{1}=N_{2}=N$, $\partial\Sigma_{i,1}=\partial\Sigma_{i,2}$, $q_{i,1}=q_{i,2}$ and $\eta_{i,1}=\eta_{i,2}$ for $i=1,\dots,N$,
Remark 3.4.
In Theorems 3.2 and 3.3, it is assumed that the coefficients of the reflective index follow the form specified in (3.1) and are real, while the conductive boundary parameter is a real constant. This assumption is also crucial in Lemma 3.2, as the proof of Theorems 3.2 and 3.3 relies on Lemma 3.2. Remark 3.5 suggests that the conclusion of Lemma 3.2 may hold in a more general scenario where the coefficients of the reflective index and the conductive boundary parameter are complex numbers. Hence, it is speculated that the linear polynomial reflective index with complex coefficients can be uniquely determined through a single far-field measurement.
Consider an admissible conductive polygonal-nest or polygonal-cell medium scatterer $(\Omega;q,\eta)$ as defined in Definition 3.5, where $\Omega\Subset\mathbb{R}^{2}$ possesses a polygonal-nest or polygonal-cell structure. Each vertex $\mathbf{x}_{c}$ of $\Omega$ corresponds to a corner with an associated opening angle $\theta\in(0,\pi)$. Utilizing the translation transformation ${\mathbf{y}}=\mathbf{x}-\mathbf{x}_{c}$, we observe that the underlying corner can be represented by (1.2), where $\Gamma_{r_{0}}^{\pm}$ denote two line segment boundaries of $\mathcal{S}_{r_{0}}$. Recall that $\vartheta_{m}$ and $\vartheta_{M}$ are the polar angles of $\Gamma_{r_{0}}^{-}$ and $\Gamma_{r_{0}}^{+}$, respectively, such that $\theta=\vartheta_{M}-\vartheta_{m}$.
Furthermore, we introduce the rotation transformation $\widetilde{\mathbf{x}}=M(\vartheta_{m})\mathbf{y}$, employing the rotation matrix $M(\vartheta_{m})$ given by:
$M(\vartheta_{m})=\begin{bmatrix}\cos(\vartheta_{m})&\sin(\vartheta_{m})\\ -\sin(\vartheta_{m})&\cos(\vartheta_{m})\end{bmatrix}.$ (3.29)
In this new coordinate system $\mathcal{C}_{N}$, we set $\mathbf{x}_{c}=\mathbf{0}$ and $\vartheta_{m}=0$. Proposition 3.1 demonstrates that if two linear polynomials coincide in the new coordinate system $\mathcal{C}_{N}$, they are identical in the original coordinate system. This result is crucial for proving the unique determination of a linear polynomial refractive index for $(\Omega;q,\eta)$. Similarly, Proposition 3.2 investigates the corresponding admissible conditions in the new coordinate system $\mathcal{C}_{N}$ for the admissible conditions (3.8) and (3.9), originally formulated in the old coordinate system. The proofs of Proposition 3.1 and 3.2 are omitted.
Proposition 3.1.
In a fixed coordinate system $\mathcal{C}$, consider two linear polynomials with the form given by
$\alpha(\mathbf{x})=a_{0}+\mathbf{a}^{\top}\mathbf{x},\ \mbox{and}\ \gamma(\mathbf{x})=b_{0}+\mathbf{b}^{\top}\mathbf{x},$
where $\mathbf{a}=(a_{1},a_{2})^{\top}$, $\mathbf{b}=(b_{1},b_{2})^{\top}$ and $\mathbf{x}=(x_{1},x_{2})^{\top}$. Pick up any fixed point $\mathbf{x}_{c}$ in the coordinate system $\mathcal{C}$ and a fixed rotation matrix $M(\vartheta_{m})$ given by (3.29). Here $\vartheta_{m}\in[-\pi,\pi)$ is a fixed angle in the coordinate system $\mathcal{C}$. Let
$\tilde{\mathbf{x}}=M(\vartheta_{m})(\mathbf{x}-\mathbf{x}_{c}),$
and
$p(\tilde{\mathbf{x}})=\alpha(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+\mathbf{x}_{c}),\ g(\tilde{\mathbf{x}})=\gamma(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+\mathbf{x}_{c}).$
If $p(\tilde{\mathbf{x}})\equiv g(\tilde{\mathbf{x}}),$ then $\alpha(\mathbf{x})=\gamma(\mathbf{x}).$
Proposition 3.2.
Consider a fixed coordinate system $\mathcal{C}$ and an admissible conductive polygonal-nest or polygonal-cell medium scatterer $(\Omega;q,\eta)$ in $\mathcal{C}$ as defined in Definition 3.5, where $\Omega$ possesses a polygonal-nest or polygonal-cell structure. Each vertex $\mathbf{x}_{c}$ of $\Omega$ corresponds to a corner with an associated opening angle $\theta\in(0,\pi)$. Let $u$ represent the total wave field of the direct scattering problem (1.7) associated with $(\Omega;q,\eta)$. Define $U(\tilde{\mathbf{x}})=u(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+{\mathbf{x}_{c}})$, where $\vartheta_{m}\in[-\pi,\pi)$ is a fixed angle in the coordinate system $\mathcal{C}$. Then, we have the following relations:
  • 1.
    If $u(\mathbf{x}_{c})=0$, then $U(\mathbf{0})=0$.
  • 2.
    If $\nabla u(\mathbf{x}_{c})\neq 0$ and $u(\mathbf{x}_{c})=0$, then $\nabla U(\mathbf{0})\neq 0$ and $U(\mathbf{0})=0$.
@@
Recall that $\mathcal{S}_{r_{0}},\Gamma_{r_{0}}^{\pm}$ and $B_{r_{0}}$ are defined in (1.2). Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDE system:
@@
where $\eta_{1}$ and $\eta_{2}$ are real constants and $q_{1},\ q_{2},\ q_{3}$ with forms
@@
where $q_{j}^{(0)}$, $q_{j}^{(1)}$ and $q_{j}^{(2)}$ are real constants, $j=1,2,3$.
Assume that for $i=1$ or $2$, either
$|u_{i}(\mathbf{0})|\neq 0,$ (3.31)
or
$\nabla u_{i}(\mathbf{0})\neq 0,\ u_{i}(\mathbf{0})=0$ (3.32)
is fulfilled. Additionally, if (3.32) holds, it is further required that $\vartheta_{M}-\vartheta_{m}\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\vartheta_{M}-\vartheta_{m}$ is the opening angle of $\mathcal{S}_{r_{0}}$.
@@
Since $u^{+}_{1}=u_{2}^{+}$ in $B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}$, we have that $u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}$ by the trace theorem, which indicates that $u_{1}^{-}=u_{1}^{+}=u_{2}^{+}=u_{2}^{-}\mbox{on}\ \Gamma_{r_{0}}^{\pm}$ according to the transmission boundary condition. Without of loss of generality, we assume that $u_{2}^{-}$ fulfills (3.31) or (3.32). Denote $\tilde{u}:=u^{-}_{1}-u_{2}^{-}$, then it can be shown directly that
@@
According to Lemma 3.1, we have $u_{2}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$. Therefore, it follows that
$u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-},\ |\delta u_{2}^{-}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{-}\|_{C^{1,\alpha}}.$ (3.34)
For convenience, we denote $\tilde{\delta}u_{2}^{-}=\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-}.$
First, we are going to prove $\eta_{1}=\eta_{2}$ and $q_{1}=q_{2}$. According to the admissible conditions (3.8) and (3.9), in the rest of this proof, we will consider two distinct cases.
Case (I): $u_{2}^{-}(\mathbf{0})\not=0.$ Using (3.33) and (3.34), we can directly see that
@@
$I_{3}$ $=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}\tilde{\delta}u_{2}^{-}u_{0}(\mathbf{x})\mathrm{d}\sigma,I_{4}=\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$
@@
Next, using $\eta_{1}=\eta_{2}$, the integral identity (3.35) can be rewritten as
@@
$\mathcal{P}_{1}$ $=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x},\ \mathcal{P}_{2}=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$

### v3

@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
$|\eta_{1}u_{1}^{-}(\mathbf{0})|=0.$
However, this contradicts to the admissible condition $u_{1}^{-}(\mathbf{0})\neq 0$ due to $\eta_{1}\neq 0$.
Case (b): $\partial_{1}u_{1}^{-}(\mathbf{0})\not=0$, $u_{1}^{-}(\mathbf{0})=\partial_{2}u_{1}^{-}(\mathbf{0})=0$. Similar to (3.15), we have the following integral equality:
$\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma=-\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma$
$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{1}^{-}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.19)
@@
$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})|}{\tau^{2}}\left|\frac{\cos\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\cos\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau}),$ (3.21)
@@
$|\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})|=0,$
which implies that $\partial_{1}u_{1}^{-}(\mathbf{0})=0$ by noting $\eta_{1}\not=0$. Hence, we obtain a contradiction.
Case (c): $\partial_{2}u_{1}^{-}(\mathbf{0})\not=0,$ $u_{1}^{-}(\mathbf{0})=\partial_{1}u_{1}^{-}(\mathbf{0})=0$. Similar to (3.15), it holds that
$\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma=-\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma$
$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{1}^{-}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.22)
@@
$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})|}{\tau^{2}}\left|\frac{\sin\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\sin\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau})$ (3.23)
@@
$|\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})|=0,$
Hence, it yields that $\partial_{2}u_{1}^{-}(\mathbf{0})=0$ since $\eta_{1}\neq 0$, which contradicts to $\partial_{2}u_{1}^{-}(\mathbf{0})\not=0$.
Case (d): $\partial_{1}u_{1}^{-}(\mathbf{0})\not=0$,$\partial_{2}u_{1}^{-}(\mathbf{0})\not=0$, $u_{1}^{-}(\mathbf{0})=0$. Similar to (3.15), we have
$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}$ $(\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\sigma-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{1}^{-}u_{0}\mathrm{d}\sigma$
$-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\psi(\mathbf{x})\mathrm{d}\sigma+I_{4}.$ (3.24)
@@
$\mathcal{N}=$ $(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{M})e^{2\mathrm{i}\vartheta_{m}}$
$\quad+(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{m})e^{2\mathrm{i}\vartheta_{M}},$ (3.25)
@@
$\mathcal{N}=$ $(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{M})e^{-2\mathrm{i}\vartheta_{m}}$
$\quad+(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{m})e^{-2\mathrm{i}\vartheta_{M}}.$ (3.26)
Let us consider the real and imaginary parts of $\partial_{1}u_{1}^{-}(\mathbf{0})$ and $\partial_{2}u_{1}^{-}(\mathbf{0})$, seperately. Denote $\partial_{1}u_{1}^{-}(\mathbf{0})=a_{1}+b_{1}\mathrm{i}$ and $\partial_{2}u_{1}^{-}(\mathbf{0})=a_{2}+b_{2}\mathrm{i}$, where $a_{i},b_{i}\in\mathbb{R}\ (i=1,2).$ For (3.25), it follows that
@@
Therefore, by (3.29), we have $a_{i}=b_{i}=0$ for $i=1,2$, which indicates that $\partial_{1}u_{1}^{-}(\mathbf{0})=\partial_{2}u_{1}^{-}(\mathbf{0})=0$. This leads to a contradiction.
@@
The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$, where $q_{i,\ell}$ follows the form specified in (3.1), while $\tilde{\eta}_{\ell}$ $\in\mathbb{C}$. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega;q_{\ell},\eta_{\ell})$. Assume that
@@
The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$ , where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ $\in\mathbb{C}$. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega_{\ell};q_{\ell},\eta_{\ell})$. Assume that
@@
for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$. Then we have $N_{1}=N_{2}=N$, $\partial\Sigma_{i,1}=\partial\Sigma_{i,2}$, $q_{i,1}=q_{i,2}$ and $\eta_{i,1}=\eta_{i,2}$ for $i=1,\dots,N$.
@@
Recall that $\mathcal{S}_{r_{0}},\Gamma_{r_{0}}^{\pm}$ and $B_{r_{0}}$ are defined in (1.2). Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDEs system:
@@
where $\eta_{1}$, $\eta_{2}$ $\in\mathbb{C}$ and $q_{1},\ q_{2},\ q_{3}$ with forms
@@
where $q_{j}^{(0)}$, $q_{j}^{(1)}$ and $q_{j}^{(2)}$ $\in\mathbb{C}$, $j=1,2,3$.
Assume that
$\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{0},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u_{i}(\mathbf{0})|\mathrm{d}\mathbf{x}\neq 0,\ i=1\ \mbox{or}\ 2.$ (3.31)
@@
Since $u^{+}_{1}=u_{2}^{+}$ in $B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}$, we have that $u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}$ by the trace theorem, which indicates that $u_{1}^{-}=u_{1}^{+}=u_{2}^{+}=u_{2}^{-}\mbox{on}\ \Gamma_{r_{0}}^{\pm}$ according to the transmission boundary condition. Without loss of generality, we assume that $u_{2}^{-}$ fulfills (3.31). Denote $\tilde{u}:=u^{-}_{1}-u_{2}^{-}$, then it can be shown directly that
@@
First, we demonstrate that the solution $u_{i}^{-}\in H_{1}(\mathcal{S}_{r_{0}}),i=1,2$ of the system (3.30) belong to $C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$. From [39, 43], we know that $u_{i}^{-}$ admits a decomposition into a regular part and a singular part in the following form:
$u_{i}^{-}=u_{i,reg}^{-}+u_{i,sing}^{-}.$
The regular part $u_{i,reg}^{-}\in H^{2}(\mathcal{S}_{r_{0}})$ and hence by the Sobolev embedding theorem, it follows that $u_{i,reg}^{-}\in C^{\beta}(\mathcal{S}_{r_{0}}),\beta\in(0,1]$. Moreover, $u_{{i,sing}}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$ is the Hölder exponent depending on the opening angle of $\mathcal{S}_{r_{0}}$. Therefore, $u_{i}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}}).$ Thus, $u_{2}^{-}$ satisfies the following expansion:
$u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\delta u_{2}^{-},\ |\delta u_{2}^{-}|\leq|\mathbf{x}|^{\alpha}\|u_{2}^{-}\|_{C^{\alpha}}.$ (3.33)
Now, we are going to prove that $\eta_{1}=\eta_{2}$ and $q_{1}=q_{2}$ in two parts.
Part 1: $\eta_{1}=\eta_{2}$. Using (3.32) and (3.33), we can directly see that
@@
$I_{3}$ $=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{-}u_{0}(\mathbf{x})\mathrm{d}\sigma,I_{4}=\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$
@@
Part 2: $q_{1}=q_{2}$. Using $\eta_{1}=\eta_{2}$, the integral identity (3.34) can be rewritten as
@@
$\mathcal{P}_{1}$ $=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x},\ \mathcal{P}_{2}=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$